\begin{thebibliography}{99}
\bibitem{2800-belle}R.~Mizuk {\it et al.}, (Belle Collaboration),
Phys. Rev. Lett. {\bf94}, 122002 (2005).


\bibitem{2800-babar}B.~Aubert {\it et al.}, (BABAR Collaboration), Phys. Rev. D {\bf78}, 112003 (2008).


\bibitem{2940-babar}B.~Aubert {\it et al.}, (BABAR Collaboration), Phys. Rev. Lett. {\bf98},
012001 (2007).


\bibitem{2940-belle}R.~Mizuk {\it et al.}, (Belle Collaboration), Phys. Rev. Lett. {\bf98}, 262001
(2007).



\bibitem{quark-model}S.~Capstick and N.~Isgur, Phys. Rev. D {\bf34}, 2809 (1986).


\bibitem{Ebert}D.~Ebert, R.~N.~Faustova, and V.~O.~Galkin, Phys. Lett. B {\bf659}, 612 (2008).



\bibitem{Cheng}H.~Y.~Cheng and C.~K.~Chua, Phys. Rev. D {\bf75}, 014006 (2007).




\bibitem{Chen}C.~Chen, X.~L.~Chen, X.~Liu, W.~Z.~Deng, and S.~L.~Zhu, Phys.
Rev. D {\bf75}, 094017 (2007).



\bibitem{Zhong}X.~H.~Zhong and Q.~Zhao, Phys. Rev. D {\bf77}, 074008 (2008).



\bibitem{Garcilazo}H.~Garcilazo, J.~Vijande, and A.~Valcarce,
J. Phys. G: Nucl. Part. Phys. {\bf34}, 961 (2007).



\bibitem{BChen}B.~Chen, D.~X.~Wang, and A.~Zhang, Chin. Phys. C {\bf33}, 1327
(2009).

\bibitem{JHe}J.~He, Z.~Ouyang, X.~Liu, and X.~Q.~Li, Phys. Rev. D {\bf84}, 114010 (2011).



\bibitem{Lutz}M.~F.~M.~Lutz and E.~E.~Kolomeitsev, Nucl. Phys. A {\bf755}, 29c (2005).


\bibitem{Tejero}C.~E.~Jim\'{e}nez-Tejero, A.~Ramos, and I.~Vida\~{n}a, Phys. Rev. C {\bf80}, 055206 (2009).

\bibitem{Dong}Y.~Dong, A.~Faessler, T.~Gutsche, and V.~E.~Lyubovitskij,
Phys. Rev. D {\bf81}, 074011 (2010); Phys.
Rev. D {\bf81}, 014006 (2010).



\bibitem{He}X.~G.~He, X.~Q.~Li, X.~Liu, and X.~Q.~Zeng, Eur. Phys. J. C {\bf51},
883 (2007).





\bibitem{Dong1}Y.~Dong, A.~Faessler, T.~Gutsche, S.~Kumano, and
V.~E.~Lyubovitskij, Phys. Rev. D {\bf82}, 034035 (2010); Phys. Rev. D {\bf83}, 094005 (2011).




\bibitem{JHe2}J.~He and X.~Liu, Phys. Rev. D {\bf82}, 114029 (2010).


\bibitem{Recio}C.~Garc\'{\i}a-Recio {\it et al.}, Phys. Rev. D {\bf79}, 054004 (2009).


\bibitem{Ortega}P.~G.~Ortega, D.~R.~Entem, and F.~Fern\'{a}ndez, arXiv:1210.2633 [hep-ph].

\bibitem{ZJR}J.~R.~Zhang, arXiv:1212.5325 [hep-ph], to appear in Phys. Rev. D (2014).

\bibitem{svzsum}M.~A.~Shifman, A.~I.~Vainshtein, and V.~I.~Zakharov, Nucl. Phys. {\bf B147}, 385 (1979); {\bf B147}, 448 (1979);
 V.~A.~Novikov, M.~A.~Shifman, A.~I.~Vainshtein, and V.~I.~Zakharov, Fortschr. Phys. {\bf 32}, 585 (1984).




\bibitem{overview1}B.~L.~Ioffe, in The Spin Structure of The Nucleon, edited by
B.~Frois, V.~W.~Hughes, and N.~de Groot (World Scientific,
Singapore, 1997).




\bibitem{overview2}S.~Narison, Camb. Monogr. Part. Phys. Nucl. Phys. Cosmol. {\bf17}, 1
(2002).




\bibitem{overview3}P.~Colangelo and A.~Khodjamirian, in At the Frontier of
Particle Physics: Handbook of QCD, edited by M.~Shifman,
Boris Ioffe Festschrift Vol. 3 (World Scientific,
Singapore, 2001), pp. 1495-1576; A. Khodjamirian,
Continuous Advances in QCD 2002/ARKADYFEST, arXiv:0209166 [hep-ph].


\bibitem{overview4}M.~Neubert, Phys. Rev. D {\bf 45},
2451 (1992); Phys. Rep. {\bf 245}, 259 (1994).


\bibitem{XYZ}M.~Nielsen, F.~S.~Navarra, and S.~H.~Lee, Phys. Rep. {\bf497},
41 (2010).



\bibitem{pentaquark}S.~L.~Zhu, Phys. Rev. Lett {\bf 91},
232002 (2003); W.~Wei, P.~Z.~Huang, H.~X.~Chen, and S.~L.~Zhu, JHEP {\bf07}, 015 (2005).



\bibitem{pentaquark1}R.~D.~Matheus, F.~S.~Navarra, M.~Nielsen, R.~Rodrigues da
Silva, and S.~H.~Lee, Phys. Lett. B {\bf578}, 323 (2004);
J.~Sugiyama, T.~Doi, and M.~Oka, Phys. Lett. B {\bf581}, 167
(2004);
M.~Eidemuller, Phys. Lett. B {\bf597}, 314 (2004);
Z.~G.~Wang, W.~M.~Yang, and S.~L.~Wan, Phys. Rev. D {\bf 72},
034012 (2005).



\bibitem{pentaquark-heavy}H.~Kim, S.~H.~Lee, and Y.~Oh, Phys. Lett. B {\bf595}, 293
(2004).



\bibitem{reinders}L.~J.~Reinders, H.~R.~Rubinstein, and S.~Yazaki, Phys. Rep. {\bf 127}, 1 (1985).

\bibitem{baryon-current}B.~L.~Ioffe, Nucl. Phys. {\bf B188}, 317 (1981);
E.~V.~Shuryak, Nucl. Phys. {\bf B198}, 83 (1982).

\bibitem{Zhang}J.~R.~Zhang and M.~Q.~Huang, Phys. Rev. D {\bf77}, 094002 (2008); Phys. Rev. D
{\bf78}, 094007 (2008); Phys. Lett. B {\bf674}, 28 (2009).

\bibitem{Zhang1}J.~R.~Zhang and M.~Q.~Huang, Phys. Rev. D {\bf80}, 056004 (2009);
J. Phys. G: Nucl. Part. Phys. {\bf37}, 025005 (2010); Commun. Theor. Phys. {\bf54}, 1075 (2010);
J.~R.~Zhang, M.~Zhong, and M.~Q.~Huang, Phys. Lett. B {\bf704}, 312 (2011).


\bibitem{HXChen}H.~X.~Chen, A. Hosaka, and S. L. Zhu, Phys. Lett. B {\bf650},
369 (2007).

\bibitem{ZGWang}Z.~G.~Wang, Nucl. Phys. A {\bf791}, 106 (2007).



\bibitem{Matheus}R.~D.~Matheus, F.~S.~Navarra, M.~Nielsen, and R.~Rodrigues da Silva, Phys. Rev. D {\bf76}, 056005 (2007).


\bibitem{Zs}J.~R.~Zhang, L.~F.~Gan, and M.~Q.~Huang, Phys. Rev. D {\bf85}, 116007 (2012);
J.~R.~Zhang and G.~F.~Chen, Phys. Rev. D {\bf86}, 116006 (2012);
J.~R.~Zhang, Phys. Rev. D {\bf87}, 076008 (2013).
\end{thebibliography}

\documentclass[showpacs,aps,prd]{revtex4}
\input epsf

\begin{document}

\title{Possibility of $\Sigma_{c}(2800)$ and $\Lambda_{c}(2940)^{+}$ as $S$-wave $D^{(*)}N$ molecular states}
\author{Jian-Rong Zhang}
\affiliation{Department of Physics, College of Science, National University of Defense Technology,
Changsha 410073, Hunan, People's Republic of China}

\begin{abstract}
In this talk, we
give a brief review of our work
studying the possibility of $\Sigma_{c}(2800)$ and $\Lambda_{c}(2940)^{+}$ as $S$-wave $D^{(*)}N$ molecular states
from QCD sum rules.
\end{abstract}
\pacs {11.55.Hx, 12.38.Lg, 12.39.Mk}\maketitle

In 2005, Belle Collaboration observed
an isotriplet of new states $\Sigma_{c}(2800)$ \cite{2800-belle},
the neutral one of which was confirmed by Babar Collaboration  \cite{2800-babar}.
Moreover, Babar collaboration reported
the observation of $\Lambda_{c}(2940)^{+}$ \cite{2940-babar},
which was confirmed by Belle Collaboration \cite{2940-belle}.

There have appeared two different ways to understand the inner structures of $\Sigma_{c}(2800)$ and $\Lambda_{c}(2940)^{+}$.
From a potential model prediction,
their masses
are close to theoretical values of $\Sigma_{c}^{*}$ with $J^{P}=\frac{3}{2}^{-}$ or $\frac{5}{2}^{-}$ and $\Lambda_{c}^{*}$ with $J^{P}=\frac{5}{2}^{-}$ or $\frac{3}{2}^{+}$, respectively \cite{quark-model}.
One way of studies grounds on the
assignments of them as conventional charmed baryons \cite{Ebert,Cheng,Chen,Zhong,Garcilazo,BChen,JHe}.
Another way bases on the assumption that they are some molecular candidates \cite{Lutz,Tejero,Dong,He,Dong1,JHe2,Recio}.
In particular, Ortega {\it et al.} studied $\Lambda_{c}(2940)^{+}$ as a $D^{*}N$ molecule
with $J^{P}=\frac{3}{2}^{-}$ in a constituent quark model, and claimed
obtaining a mass which agrees with the experimental data \cite{Ortega}.

To investigate the possibility of $\Sigma_{c}(2800)$ and $\Lambda_{c}(2940)^{+}$
as the $S$-wave $DN$ state with $J^{P}=\frac{1}{2}^{-}$ and the $S$-wave $D^{*}N$ state with $J^{P}=\frac{3}{2}^{-}$
respectively,
we devoted to studying them in Ref. \cite{ZJR} employing the method of
QCD sum rules \cite{svzsum} (for reviews
see \cite{overview1,overview2,overview3,overview4,XYZ} and references
therein; Particularly,
many theorists began to study
light pentaquark states in \cite{pentaquark,pentaquark1} and heavy ones in
\cite{pentaquark-heavy}).

The basic point of QCD sum rules is to construct a proper
interpolating current to represent the studied state.
At present, molecular currents are built up with the color-singlet currents of
composed hadrons to form hadron-hadron
configurations of fields.
One could find meson currents in Ref. \cite{reinders} and
nucleon ones in \cite{baryon-current}.
Therefore, we build following forms of currents:
\begin{eqnarray}
j&=&(\bar{q}^{c^{'}}i\gamma_{5}Q^{c^{'}})(\varepsilon_{abc}q_{1}^{Ta}C\gamma_{\mu}q_{2}^{b}\gamma_{5}\gamma^{\mu}q_{3}^{c}),
\end{eqnarray}
for the $S$-wave $DN$ or $\bar{B}N$ molecular state with $J^{P}=\frac{1}{2}^{-}$, and
\begin{eqnarray}
j^{\rho}&=&(\bar{q}^{c^{'}}\gamma^{\rho}Q^{c^{'}})(\varepsilon_{abc}q_{1}^{Ta}C\gamma_{\mu}q_{2}^{b}\gamma_{5}\gamma^{\mu}q_{3}^{c}),
\end{eqnarray}
for the $S$-wave
 $D^{*}N$ or $\bar{B}^{*}N$ molecular state with $J^{P}=\frac{3}{2}^{-}$.
Here $Q$ is heavy quark $c$ or $b$, and $q_{1}$, $q_{2}$, and $q_{3}$ denote light quarks $u$ and/or $d$.
The index $T$ means matrix
transposition, $C$ is the charge conjugation matrix,
with $a$, $b$, $c$ and $c'$ as color indices.

We derive QCD
sum rules for molecular states with similar techniques
as our previous works on heavy baryons \cite{Zhang} and molecular states \cite{Zhang1}.
Ultimately,
one can gain mass sum rules \cite{ZJR}
\begin{eqnarray}\label{sumrule1}
M_{H}^{2}&=&\bigg\{\int_{m_{Q}^{2}}^{s_{0}}ds\rho_{1}(s)s
e^{-s/M^{2}}+d/d(-\frac{1}{M^{2}})\hat{B}\Pi_{1}^{\mbox{cond}}\bigg\}/
\bigg\{\int_{m_{Q}^{2}}^{s_{0}}ds\rho_{1}(s)e^{-s/M^{2}}
+\hat{B}\Pi_{1}^{\mbox{cond}}\bigg\},
\end{eqnarray}
where $M_{H}$ is the mass of the hadronic resonance, $s_0$ is the threshold parameter,
 and $M^2$ indicates the Borel parameter.
The spectral density $\rho_{1}(s)$ is given by the correlator's imaginary part.
In its calculation, one works at leading order in $\alpha_{s}$ and considers condensates up
to dimension $12$.
To keep the heavy-quark mass finite, one can use the
momentum-space expression for the heavy-quark propagator \cite{reinders}.
The light-quark part of the
correlator can be calculated in the coordinate space employing the light-quark
propagator,
which is then
Fourier-transformed to the momentum space in $D$ dimension.
The
resulting light-quark part is combined with the heavy-quark part
before it is dimensionally regularized at $D=4$.

Numerically, one could take input parameters as
$m_{c}=1.23\pm0.05~\mbox{GeV}$, $m_{b}=4.24\pm0.06~\mbox{GeV}$,
$\langle\bar{q}q\rangle=-(0.23\pm0.03)^{3}~\mbox{GeV}^{3}$, $\langle
g\bar{q}\sigma\cdot G q\rangle=m_{0}^{2}~\langle\bar{q}q\rangle$,
$m_{0}^{2}=0.8\pm0.1~\mbox{GeV}^{2}$, $\langle
g^{2}G^{2}\rangle=0.88~\mbox{GeV}^{4}$, and $\langle
g^{3}G^{3}\rangle=0.045~\mbox{GeV}^{6}$
\cite{overview2}.
For the present case, it may have some difficulty to find a standard work window critically satisfying all
the conventional rules, which has been discussed in detail for some other cases \cite{HXChen,ZGWang,Matheus,Zs}.
By releasing the rigid OPE convergence criterion, one could choose some transition range as a compromise Borel window
and arrive at
$3.75\pm0.14\pm0.08~\mbox{GeV}$
for the $S$-wave $DN$ state with $J^{P}=\frac{1}{2}^{-}$.
To investigate the effect of different factorization of $\langle q\bar{q}q\bar{q}\rangle=\kappa\langle\bar{q}q\rangle\langle\bar{q}q\rangle$, we average three results
for $\kappa=1\sim3$ and arrive at the mass value $3.64\pm0.33~\mbox{GeV}$ concisely for the
$S$-wave $DN$ state with $J^{P}=\frac{1}{2}^{-}$,
which is somewhat higher than the experimental value of $\Sigma_{c}(2800)$ even considering the result's uncertainty.
For the $S$-wave $\bar{B}N$ state with $J^{P}=\frac{1}{2}^{-}$,
one can obtain the mass value $6.97\pm0.34~\mbox{GeV}$.
Similarly, the final result is
$3.73\pm0.35~\mbox{GeV}$ for the
$S$-wave $D^{*}N$ state with $J^{P}=\frac{3}{2}^{-}$,
which is greater than the experimental data of $\Lambda_{c}(2940)^{+}$
even taking into account the uncertainty.
For the mass of $S$-wave $\bar{B}^{*}N$ state with $J^{P}=\frac{3}{2}^{-}$,
one could gain
$6.98\pm0.34~\mbox{GeV}$.

In summary, we have studied the possibility of $\Sigma_{c}(2800)$ and $\Lambda_{c}(2940)^{+}$
as $S$-wave $D^{(*)}N$ states using QCD sum rules in Ref. \cite{ZJR}. The final results are respectively greater than the experimental data of $\Sigma_{c}(2800)$ and $\Lambda_{c}(2940)^{+}$.
In view of that corresponding molecular currents are constructed from local operators
of hadrons, the possibility of $\Sigma_c(2800)$ and $\Lambda_{c}(2940)^{+}$ as molecular states
can not be arbitrarily excluded merely from these disagreements
between molecular masses using local currents and experimental data.
However, one can infer that $\Sigma_{c}(2800)$ and $\Lambda_{c}(2940)^{+}$ could not be compact states from these results.
This may suggest a limitation of the QCD sum rule
using the local current to determine whether some state is a molecular state or not.
As byproducts, masses for their
bottom partners are predicted to be $6.97\pm0.34~\mbox{GeV}$ for the $S$-wave $\bar{B}N$ state of $J^{P}=\frac{1}{2}^{-}$ and
$6.98\pm0.34~\mbox{GeV}$ for the $S$-wave $\bar{B}^{*}N$ state of $J^{P}=\frac{3}{2}^{-}$.
\begin{acknowledgments}
This work was supported by the National Natural Science
Foundation of China under Contract Nos. 11105223, 10947016, 11275268, the
Foundation of NUDT (No. JC11-02-12), and the
project in NUDT for excellent youth talents.
\end{acknowledgments}

\end{document}